OpenStudy (anonymous):

Can anyone please give me an example of an expnent with a zero power, anti powers (negative exponents), in numerator, in denominator

OpenStudy (freckles):

\[5^0=1\]

OpenStudy (rulnick):

2^-2 = (1/2)^2 = 1/4 2^-1 = 1/2 2^0 = 1 2^1 = 2

OpenStudy (rulnick):

1/(2^-1) = 2^1 = 2

OpenStudy (rulnick):

x^3 y^-2 = x^3 / y^2

OpenStudy (anonymous):

where is the anti powers ??

OpenStudy (freckles):

i don't know what antipowers are :(

OpenStudy (anonymous):

Negative exponents

OpenStudy (rulnick):

x^3 / y^-2 = x^3 y^2

OpenStudy (anonymous):